% Generated by IEEEtranN.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{54}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtranN.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem[Dinh et~al.(2001)Dinh, Hoang, and Nguyen]{diend1}
D.~Dinh, K.~Hoang, and V.-T. Nguyen, ``{Vietnamese Word Segmentation},'' in
  \emph{{Proceedings of the Sixth Natural Language Processing Pacific Rim
  Symposium}}, 2001, pp. 749--756.

\bibitem[Le et~al.(2010)Le, Roussanaly, Nguyen, and Rossignol]{lehongtag10}
H.-P. Le, A.~Roussanaly, T.-M.-H. Nguyen, and M.~Rossignol, ``{An empirical
  study of maximum entropy approach for part-of-speech tagging of Vietnamese
  texts},'' in \emph{{Traitement Automatique des Langues Naturelles - TALN
  2010}}.\hskip 1em plus 0.5em minus 0.4em\relax Montr{\'e}al, Canada: {ATALA
  (Association pour le Traitement Automatique des Langues)}, 2010, p.~12.

\bibitem[Nguyen et~al.(2009)Nguyen, Vu, Nguyen, Nguyen, and
  Le]{nguyen-etal-2009-building}
P.-T. Nguyen, X.-L. Vu, T.-M.-H. Nguyen, V.-H. Nguyen, and H.-P. Le,
  ``{Building a Large Syntactically-Annotated Corpus of {V}ietnamese},'' in
  \emph{{Proceedings of the Third Linguistic Annotation Workshop ({LAW}
  {III})}}, 2009, pp. 182--185.

\bibitem[Nguyen et~al.(2018{\natexlab{a}})Nguyen, Miyao, Le, and
  Nguyen]{10.1007/s10579-017-9398-3}
Q.~T. Nguyen, Y.~Miyao, H.~T. Le, and N.~T. Nguyen, ``{Ensuring Annotation
  Consistency and Accuracy for Vietnamese Treebank},'' \emph{{Language
  Resources and Evaluation}}, vol.~52, no.~1, p. 269–315, 2018.

\bibitem[Nguyen et~al.(2014)Nguyen, Nguyen, Pham, Nguyen, and
  Le~Nguyen]{10.1007/978-3-319-07983-7_26}
D.~Q. Nguyen, D.~Q. Nguyen, S.~B. Pham, P.-T. Nguyen, and M.~Le~Nguyen, ``{From
  Treebank Conversion to Automatic Dependency Parsing for Vietnamese},'' in
  \emph{Natural Language Processing and Information Systems}, E.~M{\'e}tais,
  M.~Roche, and M.~Teisseire, Eds.\hskip 1em plus 0.5em minus 0.4em\relax Cham:
  Springer International Publishing, 2014, pp. 196--207.

\bibitem[Thi et~al.(2013)Thi, My, Nguyen~Viet, Nguyen Thi~Minh, and
  Le~Hong]{6719884}
L.~N. Thi, L.~H. My, H.~Nguyen~Viet, H.~Nguyen Thi~Minh, and P.~Le~Hong,
  ``{Building a treebank for Vietnamese dependency parsing},'' in \emph{The
  2013 RIVF International Conference on Computing \& Communication Technologies
  - Research, Innovation, and Vision for Future (RIVF)}, 2013, pp. 147--151.

\bibitem[Nguyen(2018)]{nguyen-2018-bktreebank}
K.-H. Nguyen, ``{{BKT}reebank: Building a {V}ietnamese Dependency Treebank},''
  in \emph{Proceedings of the Eleventh International Conference on Language
  Resources and Evaluation ({LREC} 2018)}.\hskip 1em plus 0.5em minus
  0.4em\relax Miyazaki, Japan: European Language Resources Association (ELRA),
  May 2018.

\bibitem[Phuong et~al.(2017)Phuong, Hoang, Khoai, Huyen, Luong, and
  Hiep]{JCSCE}
L.~H. Phuong, P.~Hoang, P.~Khoai, N.~Huyen, N.~Luong, and N.~Hiep,
  ``{Vietnamese Semantic Role Labelling},'' \emph{VNU Journal of Science:
  Computer Science and Communication Engineering}, vol.~33, no.~2, 2017.

\bibitem[Dinh and Vu(2006)]{dinhdien06}
D.~Dinh and T.~Vu, ``{A Maximum Entropy Approach for Vietnamese Word
  Segmentation},'' in \emph{Proceedings of the 2006 International Conference on
  Research, Innovation and Vision for the Future}, 2006, pp. 248--253.

\bibitem[Ha(2003)]{ha2003}
L.~A. Ha, ``{A method for word segmentation in Vietnamese},'' in
  \emph{{Proceedings of the Corpus Linguistics 2003 Conference}}, 2003, pp.
  282--287.

\bibitem[Nguyen et~al.(2006)Nguyen, Nguyen, Phan, Nguyen, and
  Ha]{nguyenetal2006}
C.-T. Nguyen, T.-K. Nguyen, X.-H. Phan, L.-M. Nguyen, and Q.-T. Ha,
  ``{{V}ietnamese Word Segmentation with {CRF}s and {SVM}s: An
  Investigation},'' in \emph{{Proceedings of PACLIC}}.\hskip 1em plus 0.5em
  minus 0.4em\relax {Tsinghua University Press}, 2006, pp. 215--222.

\bibitem[Dinh et~al.(2008)Dinh, Le, Nguyen, Nguyen, Rossignol, and
  Vu]{thangdq1}
Q.-T. Dinh, H.-P. Le, T.-M.-H. Nguyen, C.-T. Nguyen, M.~Rossignol, and X.~L.
  Vu, ``Word segmentation of vietnamese texts: a comparison of approaches,'' in
  \emph{{Proceedings of the Eleventh International Conference on Language
  Resources and Evaluation (LREC 2008)}}, 2008, pp. 1933--1936.

\bibitem[Le et~al.(2008)Le, Nguyen, Roussanaly, and Ho]{hongphuong08}
H.-P. Le, T.-M.-H. Nguyen, A.~Roussanaly, and T.-V. Ho, ``{A Hybrid Approach to
  Word Segmentation of Vietnamese Texts},'' in \emph{{Language and Automata
  Theory and Applications}}.\hskip 1em plus 0.5em minus 0.4em\relax {Springer
  Berlin Heidelberg}, 2008, pp. 240--249.

\bibitem[{Song Nguyen Duc Cong} et~al.(2016){Song Nguyen Duc Cong}, {Quoc Hung
  Ngo}, and {Jiamthapthaksin}]{songnguyen16}
{Song Nguyen Duc Cong}, {Quoc Hung Ngo}, and R.~{Jiamthapthaksin},
  ``{State-of-the-art Vietnamese word segmentation},'' in \emph{{2016 2nd
  International Conference on Science in Information Technology (ICSITech)}},
  2016, pp. 119--124.

\bibitem[Nguyen and Le(2016)]{phongnt1}
T.~P. Nguyen and A.~C. Le, ``{A hybrid approach to Vietnamese word
  segmentation},'' in \emph{{Proceeding of IEEE-RIVF}}, 2016, pp. 114--119.

\bibitem[Nguyen et~al.(2018{\natexlab{b}})Nguyen, Nguyen, Vu, Dras, and
  Johnson]{datnq1}
D.~Q. Nguyen, D.~Q. Nguyen, T.~Vu, M.~Dras, and M.~Johnson, ``{A Fast and
  Accurate Vietnamese Word Segmenter},'' in \emph{{Proceedings of LREC}}, 2018,
  pp. 2582--2587.

\bibitem[Nguyen et~al.(2020{\natexlab{a}})Nguyen, Van~Thin, Van~Nguyen, and
  Nguyen]{vund2018}
D.-V. Nguyen, D.~Van~Thin, K.~Van~Nguyen, and N.~L.-T. Nguyen, ``{Vietnamese
  Word Segmentation with SVM: Ambiguity Reduction and Suffix Capture},'' in
  \emph{Computational Linguistics}, L.-M. Nguyen, X.-H. Phan, K.~Hasida, and
  S.~Tojo, Eds.\hskip 1em plus 0.5em minus 0.4em\relax Singapore: Springer
  Singapore, 2020, pp. 400--413.

\bibitem[Nguyen et~al.(2021{\natexlab{a}})Nguyen, Vo, Thin, and
  Nguyen]{spanviws}
D.-V. Nguyen, L.-B. Vo, D.~V. Thin, and N.~L.-T. Nguyen, ``{Span Labeling
  Approach for Vietnamese and Chinese Word Segmentation},'' in \emph{PRICAI
  2021: Trends in Artificial Intelligence}, D.~N. Pham, T.~Theeramunkong,
  G.~Governatori, and F.~Liu, Eds.\hskip 1em plus 0.5em minus 0.4em\relax Cham:
  Springer International Publishing, 2021, pp. 244--258.

\bibitem[{Pham} et~al.(2009){Pham}, {Tran}, and {Pham}]{ducpd1}
D.~D. {Pham}, G.~B. {Tran}, and S.~B. {Pham}, ``{A Hybrid Approach to
  Vietnamese Word Segmentation Using Part of Speech Tags},'' in \emph{{2009
  International Conference on Knowledge and Systems Engineering}}, 2009, pp.
  154--161.

\bibitem[Nguyen et~al.(2017)Nguyen, Vu, Nguyen, Dras, and
  Johnson]{nguyen-etal-2017-word}
D.~Q. Nguyen, T.~Vu, D.~Q. Nguyen, M.~Dras, and M.~Johnson, ``{From Word
  Segmentation to {POS} Tagging for {V}ietnamese},'' in \emph{{Proceedings of
  the Australasian Language Technology Association Workshop 2017}}, 2017, pp.
  108--113.

\bibitem[Nguyen(2019)]{nguyen-2019-neural}
D.~Q. Nguyen, ``{A neural joint model for {V}ietnamese word segmentation, {POS}
  tagging and dependency parsing},'' in \emph{{Proceedings of the The 17th
  Annual Workshop of the Australasian Language Technology Association}}, 2019,
  pp. 28--34.

\bibitem[Kieu and Pham(2010)]{5632131}
B.~T. Kieu and S.~B. Pham, ``{Sentiment Analysis for Vietnamese},'' in
  \emph{2010 Second International Conference on Knowledge and Systems
  Engineering}, 2010, pp. 152--157.

\bibitem[Duyen et~al.(2014)Duyen, Bach, and Phuong]{7043403}
N.~T. Duyen, N.~X. Bach, and T.~M. Phuong, ``{An empirical study on sentiment
  analysis for Vietnamese},'' in \emph{2014 International Conference on
  Advanced Technologies for Communications (ATC 2014)}, 2014, pp. 309--314.

\bibitem[Nguyen et~al.(2019)Nguyen, Nguyen, Ngo, Vu, Tran, Ngo, and
  Le]{vlsp2016sa}
H.~T.~M. Nguyen, H.~V. Nguyen, Q.~T. Ngo, L.~X. Vu, V.~M. Tran, B.~X. Ngo, and
  C.~A. Le, ``{VLSP SHARED TASK: SENTIMENT ANALYSIS},'' \emph{Journal of
  Computer Science and Cybernetics}, vol.~34, no.~4, p. 295–310, Jan. 2019.

\bibitem[Vo et~al.(2017)Vo, Nguyen, Le, and Nguyen]{8119429}
Q.-H. Vo, H.-T. Nguyen, B.~Le, and M.-L. Nguyen, ``Multi-channel lstm-cnn model
  for vietnamese sentiment analysis,'' in \emph{2017 9th International
  Conference on Knowledge and Systems Engineering (KSE)}, 2017, pp. 24--29.

\bibitem[Bach et~al.(2015)Bach, Pham, Tai, and Phuong]{7371776}
N.~X. Bach, D.~V. Pham, N.~D. Tai, and T.~M. Phuong, ``{Mining Vietnamese
  Comparative Sentences for Sentiment Analysis},'' in \emph{2015 Seventh
  International Conference on Knowledge and Systems Engineering (KSE)}, 2015,
  pp. 162--167.

\bibitem[Vo et~al.(2016)Vo, Lam, Nguyen, and Tuong]{bagofstruct}
H.~T. Vo, H.~C. Lam, D.~D. Nguyen, and N.~H. Tuong, ``{Topic classification and
  sentiment analysis for Vietnamese education survey system},'' \emph{{Asian
  Journal of Computer Science and Information Technology}}, vol.~6, no.~3, pp.
  27--34, 2016.

\bibitem[Tran and Phan(2017)]{Tran2017}
T.~K. Tran and T.~T. Phan, ``{Mining opinion targets and opinion words from
  online reviews},'' \emph{International Journal of Information Technology},
  vol.~9, no.~3, pp. 239--249, Sep 2017.

\bibitem[Ha et~al.(2011)Ha, Vu, Pham, and Luu]{10.1007/978-3-642-23620-4_21}
Q.-T. Ha, T.-T. Vu, H.-T. Pham, and C.-T. Luu, ``{An Upgrading Feature-Based
  Opinion Mining Model on Vietnamese Product Reviews},'' in \emph{Active Media
  Technology}, N.~Zhong, V.~Callaghan, A.~A. Ghorbani, and B.~Hu, Eds.\hskip
  1em plus 0.5em minus 0.4em\relax Berlin, Heidelberg: Springer Berlin
  Heidelberg, 2011, pp. 173--185.

\bibitem[Vu et~al.(2011)Vu, Pham, Luu, and Ha]{Vu2011}
T.-T. Vu, H.-T. Pham, C.-T. Luu, and Q.-T. Ha, \emph{{A Feature-Based Opinion
  Mining Model on Product Reviews in Vietnamese}}.\hskip 1em plus 0.5em minus
  0.4em\relax Berlin, Heidelberg: Springer Berlin Heidelberg, 2011, pp. 23--33.

\bibitem[Phan and Cao(2014)]{10.1145/2676585.2676606}
D.-H. Phan and T.-D. Cao, ``{Applying Skip-Gram Word Estimation and SVM-Based
  Classification for Opinion Mining Vietnamese Food Places Text Reviews},'' in
  \emph{Proceedings of the Fifth Symposium on Information and Communication
  Technology}, ser. SoICT '14.\hskip 1em plus 0.5em minus 0.4em\relax New York,
  NY, USA: Association for Computing Machinery, 2014, p. 232–239.

\bibitem[Pham et~al.(2016)Pham, Le, and Le]{7758052}
D.-H. Pham, A.-C. Le, and T.-K.-C. Le, ``{Learning Semantic Representations for
  Rating Vietnamese Comments},'' in \emph{2016 Eighth International Conference
  on Knowledge and Systems Engineering (KSE)}, 2016, pp. 193--198.

\bibitem[Nguyen et~al.(2018{\natexlab{c}})Nguyen, Nguyen, and Nguyen]{8573351}
V.~D. Nguyen, K.~V. Nguyen, and N.~L.-T. Nguyen, ``{Variants of Long Short-Term
  Memory for Sentiment Analysis on Vietnamese Students’ Feedback Corpus},''
  in \emph{2018 10th International Conference on Knowledge and Systems
  Engineering (KSE)}, 2018, pp. 306--311.

\bibitem[Truong et~al.(2020)Truong, Le, and Le-Dang]{9335912}
T.-L. Truong, H.-L. Le, and T.-P. Le-Dang, ``{Sentiment Analysis Implementing
  BERT-based Pre-trained Language Model for Vietnamese},'' in \emph{2020 7th
  NAFOSTED Conference on Information and Computer Science (NICS)}, 2020, pp.
  362--367.

\bibitem[Nguyen et~al.(2018{\natexlab{d}})Nguyen, Nguyen, Nguyen, Truong, and
  Nguyen]{vsfc}
K.~V. Nguyen, V.~D. Nguyen, P.~X.~V. Nguyen, T.~T.~H. Truong, and N.~L.-T.
  Nguyen, ``{UIT-VSFC: Vietnamese Students’ Feedback Corpus for Sentiment
  Analysis},'' in \emph{2018 10th International Conference on Knowledge and
  Systems Engineering (KSE)}, 2018, pp. 19--24.

\bibitem[Ho et~al.(2020)Ho, Nguyen, Nguyen, Pham, Nguyen, Nguyen, and
  Nguyen]{10.1007/978-981-15-6168-9_27}
V.~A. Ho, D.~H.-C. Nguyen, D.~H. Nguyen, L.~T.-V. Pham, D.-V. Nguyen, K.~V.
  Nguyen, and N.~L.-T. Nguyen, ``{Emotion Recognition for Vietnamese Social
  Media Text},'' in \emph{Computational Linguistics}, L.-M. Nguyen, X.-H. Phan,
  K.~Hasida, and S.~Tojo, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  Singapore: Springer Singapore, 2020, pp. 319--333.

\bibitem[Van~Thin et~al.(2020)Van~Thin, Nguyen, Van~Nguyen, Nguyen, and
  Nguyen]{10.1007/978-981-15-6168-9_15}
D.~Van~Thin, D.-V. Nguyen, K.~Van~Nguyen, N.~L.-T. Nguyen, and A.~H.-T. Nguyen,
  ``{Multi-task Learning for Aspect and Polarity Recognition on Vietnamese
  Datasets},'' in \emph{Computational Linguistics}, L.-M. Nguyen, X.-H. Phan,
  K.~Hasida, and S.~Tojo, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  Singapore: Springer Singapore, 2020, pp. 169--180.

\bibitem[Nguyen et~al.(2021{\natexlab{b}})Nguyen, Le, Tran, Luu, Hoang, and
  Phan]{DBLP:conf/somet/0002LTLHP21}
H.~D. Nguyen, T.~Le, K.~V. Tran, S.~T. Luu, S.~N. Hoang, and H.~T. Phan,
  ``{Multi-Level Sentiment Analysis of Product Reviews Based on Grammar
  Rules},'' in \emph{Proceedings of SoMeT 2021}, H.~Fujita and
  H.~P{\'{e}}rez{-}Meana, Eds., vol. 337.\hskip 1em plus 0.5em minus
  0.4em\relax {IOS} Press, 2021, pp. 444--456.

\bibitem[Nguyen et~al.(2020{\natexlab{b}})Nguyen, Huynh, Hoang, Pham, and
  Zelinka]{DBLP:conf/enase/0002HHPZ20}
H.~D. Nguyen, T.~Huynh, S.~N. Hoang, V.~T. Pham, and I.~Zelinka,
  ``{Language-oriented Sentiment Analysis based on the Grammar Structure and
  Improved Self-attention Network},'' in \emph{Proceedings of the 15th
  International Conference on Evaluation of Novel Approaches to Software
  Engineering, {ENASE} 2020, Prague, Czech Republic, May 5-6, 2020}, R.~Ali,
  H.~Kaindl, and L.~A. Maciaszek, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  {SCITEPRESS}, 2020, pp. 339--346.

\bibitem[Luc~Phan et~al.(2021)Luc~Phan, Huynh~Pham, Thi-Thanh~Nguyen,
  Khai~Huynh, Thi~Nguyen, Thanh~Nguyen, Van~Huynh, and
  Van~Nguyen]{10.1007/978-3-030-82147-0_53}
L.~Luc~Phan, P.~Huynh~Pham, K.~Thi-Thanh~Nguyen, S.~Khai~Huynh, T.~Thi~Nguyen,
  L.~Thanh~Nguyen, T.~Van~Huynh, and K.~Van~Nguyen, ``{SA2SL: From Aspect-Based
  Sentiment Analysis to Social Listening System for Business Intelligence},''
  in \emph{Knowledge Science, Engineering and Management}, H.~Qiu, C.~Zhang,
  Z.~Fei, M.~Qiu, and S.-Y. Kung, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  Cham: Springer International Publishing, 2021, pp. 647--658.

\bibitem[Joulin et~al.(2017)Joulin, Grave, Bojanowski, and
  Mikolov]{joulin2017bag}
A.~Joulin, E.~Grave, P.~Bojanowski, and T.~Mikolov, ``{Bag of Tricks for
  Efficient Text Classification},'' in \emph{Proceedings of the 15th Conference
  of the European Chapter of the Association for Computational Linguistics:
  Volume 2, Short Papers}.\hskip 1em plus 0.5em minus 0.4em\relax Association
  for Computational Linguistics, April 2017, pp. 427--431.

\bibitem[Lê et~al.(2020)Lê, The~Lam, Nguyen, and Thanh~Nguyen]{9140757}
N.~C. Lê, N.~The~Lam, S.~H. Nguyen, and D.~Thanh~Nguyen, ``{On Vietnamese
  Sentiment Analysis: A Transfer Learning Method},'' in \emph{2020 RIVF
  International Conference on Computing and Communication Technologies (RIVF)},
  2020, pp. 1--5.

\bibitem[Nguyen et~al.(2020{\natexlab{c}})Nguyen, Nguyen, Luong, and
  Ngo]{9335899}
Q.~T. Nguyen, T.~L. Nguyen, N.~H. Luong, and Q.~H. Ngo, ``{Fine-Tuning BERT for
  Sentiment Analysis of Vietnamese Reviews},'' in \emph{2020 7th NAFOSTED
  Conference on Information and Computer Science (NICS)}, 2020, pp. 302--307.

\bibitem[Tran and Bui(2020)]{9287650}
O.~T. Tran and V.~T. Bui, ``{A BERT-based Hierarchical Model for Vietnamese
  Aspect Based Sentiment Analysis},'' in \emph{2020 12th International
  Conference on Knowledge and Systems Engineering (KSE)}, 2020, pp. 269--274.

\bibitem[Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova]{devlin-etal-2019-bert}
J.~Devlin, M.-W. Chang, K.~Lee, and K.~Toutanova, ``{{BERT}: Pre-training of
  Deep Bidirectional Transformers for Language Understanding},'' in
  \emph{Proceedings of the 2019 Conference of the North {A}merican Chapter of
  the Association for Computational Linguistics: Human Language Technologies,
  Volume 1 (Long and Short Papers)}.\hskip 1em plus 0.5em minus 0.4em\relax
  Minneapolis, Minnesota: Association for Computational Linguistics, Jun. 2019,
  pp. 4171--4186.

\bibitem[Bojanowski et~al.(2017)Bojanowski, Grave, Joulin, and
  Mikolov]{bojanowski2017enriching}
P.~Bojanowski, E.~Grave, A.~Joulin, and T.~Mikolov, ``{Enriching Word Vectors
  with Subword Information},'' \emph{Transactions of the Association for
  Computational Linguistics}, vol.~5, pp. 135--146, 2017.

\bibitem[Nguyen and Tuan~Nguyen(2020)]{phobert}
D.~Q. Nguyen and A.~Tuan~Nguyen, ``{{P}ho{BERT}: Pre-trained language models
  for {V}ietnamese},'' in \emph{Findings of the Association for Computational
  Linguistics: EMNLP 2020}.\hskip 1em plus 0.5em minus 0.4em\relax Online:
  Association for Computational Linguistics, Nov. 2020, pp. 1037--1042.

\bibitem[Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch]{sennrich-etal-2016-neural}
R.~Sennrich, B.~Haddow, and A.~Birch, ``{Neural Machine Translation of Rare
  Words with Subword Units},'' in \emph{Proceedings of the 54th Annual Meeting
  of the Association for Computational Linguistics (Volume 1: Long
  Papers)}.\hskip 1em plus 0.5em minus 0.4em\relax Berlin, Germany: Association
  for Computational Linguistics, Aug. 2016, pp. 1715--1725.

\bibitem[Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and
  Stoyanov]{conneau-etal-2020-unsupervised}
A.~Conneau, K.~Khandelwal, N.~Goyal, V.~Chaudhary, G.~Wenzek, F.~Guzm{\'a}n,
  E.~Grave, M.~Ott, L.~Zettlemoyer, and V.~Stoyanov, ``{Unsupervised
  Cross-lingual Representation Learning at Scale},'' in \emph{Proceedings of
  the 58th Annual Meeting of the Association for Computational
  Linguistics}.\hskip 1em plus 0.5em minus 0.4em\relax Online: Association for
  Computational Linguistics, Jul. 2020, pp. 8440--8451.

\bibitem[Gu et~al.(2022)Gu, Goel, and R\'e]{s4}
A.~Gu, K.~Goel, and C.~R\'e, ``{Efficiently Modeling Long Sequences with
  Structured State Spaces},'' in \emph{The International Conference on Learning
  Representations ({ICLR})}, 2022.

\bibitem[Baevski and Auli(2018)]{DBLP:journals/corr/abs-1809-10853}
A.~Baevski and M.~Auli, ``{Adaptive Input Representations for Neural Language
  Modeling},'' \emph{CoRR}, vol. abs/1809.10853, 2018.

\bibitem[Loshchilov and Hutter(2017)]{DBLP:journals/corr/abs-1711-05101}
I.~Loshchilov and F.~Hutter, ``{Fixing Weight Decay Regularization in Adam},''
  \emph{CoRR}, vol. abs/1711.05101, 2017.

\bibitem[Pedregosa et~al.(2011)Pedregosa, Varoquaux, Gramfort, Michel, Thirion,
  Grisel, Blondel, Prettenhofer, Weiss, Dubourg, Vanderplas, Passos,
  Cournapeau, Brucher, Perrot, and Duchesnay]{scikit-learn}
F.~Pedregosa, G.~Varoquaux, A.~Gramfort, V.~Michel, B.~Thirion, O.~Grisel,
  M.~Blondel, P.~Prettenhofer, R.~Weiss, V.~Dubourg, J.~Vanderplas, A.~Passos,
  D.~Cournapeau, M.~Brucher, M.~Perrot, and E.~Duchesnay, ``{Scikit-learn:
  Machine Learning in {P}ython},'' \emph{Journal of Machine Learning Research},
  vol.~12, pp. 2825--2830, 2011.

\bibitem[Bach et~al.(2016)Bach, Linh, and Phuong]{10.1007/978-3-319-46675-0_26}
N.~X. Bach, N.~D. Linh, and T.~M. Phuong, ``{Vietnamese POS Tagging for Social
  Media Text},'' in \emph{Neural Information Processing}, A.~Hirose, S.~Ozawa,
  K.~Doya, K.~Ikeda, M.~Lee, and D.~Liu, Eds.\hskip 1em plus 0.5em minus
  0.4em\relax Cham: Springer International Publishing, 2016, pp. 233--242.

\end{thebibliography}

\documentclass[conference]{IEEEtran}
\IEEEoverridecommandlockouts

\usepackage{amsmath,amssymb,amsfonts}
\usepackage{algorithmic}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{xcolor}
\usepackage{placeins}
\usepackage{times}
\usepackage[T5]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{microtype}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{marvosym}
\usepackage{float}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{pifont}
\usepackage{siunitx}
\usepackage{latexsym}
\usepackage{etoolbox}
\usepackage{balance}

\usepackage[numbers]{natbib}
\usepackage[bookmarks,bookmarksopen,bookmarksdepth=3,bookmarksnumbered]{hyperref}
\hypersetup{colorlinks, citecolor=blue, linkcolor=blue, urlcolor=blue}

\usepackage[left=2.5cm,right=2.5cm,top=1.75cm,bottom=1.5cm]{geometry}

\PassOptionsToPackage{bookmarks=false}{hyperref}

\begin{document}

\title{Is word segmentation necessary for\\ Vietnamese sentiment classification?}

\author{\IEEEauthorblockN{\textbf{Duc-Vu Nguyen$^{\text{1, 2}}$, Ngan Luu-Thuy Nguyen$^{\text{1, 2}}$}}
\IEEEauthorblockA{$^{\text{1}}$University of Information Technology, Ho Chi Minh City, Vietnam \\
$^{\text{2}}$Vietnam National University, Ho Chi Minh City, Vietnam\\
\texttt{\{vund,ngannlt\}@uit.edu.vn}}
}

\maketitle

\begin{abstract}
To the best of our knowledge, this paper made the first attempt to answer whether word segmentation is necessary for Vietnamese sentiment classification. To do this, we presented five pre-trained monolingual S4-based language models for Vietnamese, including one model without word segmentation, and four models using \texttt{RDRsegmenter}, \texttt{uitnlp}, \texttt{pyvi}, or \texttt{underthesea} toolkits in the pre-processing data phase. According to comprehensive experimental results on two corpora, including the VLSP2016-SA corpus of technical article reviews from the news and social media and the UIT-VSFC corpus of the educational survey, we have two suggestions. Firstly, using traditional classifiers like Naive Bayes or Support Vector Machines, word segmentation maybe not be necessary for the Vietnamese sentiment classification corpus, which comes from the social domain. Secondly, word segmentation is necessary for Vietnamese sentiment classification when word segmentation is used before using the BPE method and feeding into the deep learning model. In this way, the \texttt{RDRsegmenter} is the stable toolkit for word segmentation among the \texttt{uitnlp}, \texttt{pyvi}, and \texttt{underthesea} toolkits.
\end{abstract}

\begin{IEEEkeywords}
Natural Language Processing, Vietnamese Sentiment Classification, Vietnamese Word Segmentation, Structured State Space Sequence (S4)
\end{IEEEkeywords}

\section{Introduction}
\label{sec:introduction}
Word segmentation is a fundamental problem in the field of Vietnamese natural language processing. For instance, the Vietnamese text ``hiện đại hóa đất nước'' ($\text{modernize}_\text{hiện\_đại\_hoá}$ $\text{country}_\text{đất\_nước}$), which consists of five syllables, is segmented into ``hiện\_đại\_hoá đất\_nước''. Underscores indicate the white spaces function as syllable separators, and white spaces are used for word splits. That is a challenging problem in the early stage of natural language processing in Vietnamese \cite{diend1}. Common tasks in Vietnamese syntax analysis tasks, such as part-of-speech tagging \cite{lehongtag10}, constituency parsing \cite{nguyen-etal-2009-building, 10.1007/s10579-017-9398-3}, dependency parsing \cite{10.1007/978-3-319-07983-7_26,6719884,nguyen-2018-bktreebank}, and semantic parsing \cite{JCSCE}, are required to undergo word segmentation. In these tasks, the mistake of word segmentation directly affects them. Therefore, many previous works introduced various approaches to improve Vietnamese word segmentation performance, including single word segmentation task only \cite{diend1, dinhdien06, ha2003, nguyenetal2006, thangdq1, hongphuong08, songnguyen16, phongnt1, datnq1, vund2018, spanviws} and multi-task containing word segmentation and part-of-speech tagging \cite{ducpd1, nguyen-etal-2017-word}, dependency parsing \cite{nguyen-2019-neural}. From the previous works above, word segmentation is obligatory for Vietnamese syntax analysis tasks, including part-of-speech tagging, constituency parsing, dependency parsing, and semantic parsing.

In addition to the Vietnamese syntax analysis tasks, sentiment classification attracts many researchers because we can apply it quickly to real-life applications. Firstly, \citet{5632131} introduced a system that can classify a computer product review into positive or negative. The study of \citet{5632131} is the first work that analyzes sentiment at the sentence level in Vietnamese. Another example, in the research of \citet{7043403}, they proposed a framework classifying the hotel review into one positive, neutral, or negative class. Almost all do word segmentation in data pre-processing before researching methodology to solve the sentiment classification task. The \texttt{vnTokenizer} toolkit \cite{hongphuong08} was used for pre-processing the VLSP2016-SA \cite{vlsp2016sa} and VS corpora in the work of \citet{8119429}, the hotel reviews corpus proposed by \citet{7043403}, the electronic devices corpus proposed by \citet{7371776}, the education survey corpus proposed by \citet{bagofstruct}, and the online reviews corpus proposed by \citet{Tran2017}. Another Vietnamese word segmenter proposed by \citet{ducpd1}, was used in the works of \citet{10.1007/978-3-642-23620-4_21}, \citet{5632131}, and \citet{Vu2011} about product reviews domain. After that, \citet{10.1145/2676585.2676606} used the \texttt{JVnSegmenter} toolkit \cite{nguyenetal2006} for pre-processing their food reviews corpus. In addition, \citet{7758052} apply the \texttt{UETsegmenter} \cite{phongnt1} to preproces their mobile product reviews corpus. Lastly, the works \cite{8573351, 9335912} about sentiment analysis on feedback of students \cite{vsfc} and \cite{10.1007/978-981-15-6168-9_27} about emotion recognition used the \texttt{RDRsegmenter} of \citet{datnq1} for their research. All of the above works on Vietnamese sentiment analysis used Vietnamese word segmenter published in scientific papers. Moreover, we can survey two well-known Vietnamese word segmenters, namely \texttt{pyvi}\footnote{\url{https://pypi.org/project/pyvi/}} and \texttt{underthesea}\footnote{\url{https://pypi.org/project/underthesea/}} scientifically unpublished up to now. For instances, the \texttt{pyvi} toolkit was used in the research of \citet{10.1007/978-981-15-6168-9_15} about sentiment analysis on VLSP2018-SA corpus \cite{vlsp2016sa} and research of \citet{DBLP:conf/somet/0002LTLHP21} on product reviews. Another instance, \citet{DBLP:conf/enase/0002HHPZ20} used \texttt{underthesea} toolkit for pre-processing their electronic products comments dataset. Lastly, we observed many prior studies on Vietnamese sentiment classification apply word segmentation in the pre-processing phase.

Although we can observe that the word segmentation phase is almost necessary for the Vietnamese sentiment classification problem, recently, there have been some studies on the Vietnamese sentiment classification problem without the word segmentation phase. For example, the research \cite{10.1007/978-3-030-82147-0_53} applied the fastText \cite{joulin2017bag} model for pre-processing and embedding the input data without the Vietnamese word segmentation phase. For other examples, the research \cite{9140757, 9335899, 9287650} studied the sentiment classification problem using the pre-trained multilingual language model mBERT \cite{devlin-etal-2019-bert}, which is not required Vietnamese word segmentation. To our best knowledge, the studies \cite{10.1007/978-3-030-82147-0_53,9140757, 9335899, 9287650} applied sub-words \cite{bojanowski2017enriching} method for the better handling unseen words. On the other hand, the fastText and mBERT models are made for multilingual purposes. Hence, the studies \cite{10.1007/978-3-030-82147-0_53,9140757, 9335899, 9287650} did not use the word segmentation phase for the Vietnamese sentiment classification problem. Indeed, the research \cite{9335912} used \texttt{RDRsegmenter} toolkit for data pre-processing before using the pre-trained monolingual PhoBERT model \cite{phobert}, which is made for Vietnamese and applied Byte-Pair Encoding (BPE) method \cite{sennrich-etal-2016-neural} for sub-word representations for Vietnamese.

From the above observations, we have to admit that word segmentation is crucial for Vietnamese syntax analysis tasks, including part-of-speech tagging, constituency parsing, dependency parsing, and semantic parsing. That strongly motivates many proposed Vietnamese word segmentation methods in prior studies. On the Vietnamese Treebank corpus for word segmentation \cite{nguyen-etal-2009-building}, the highest F-score is 98.31\% achieved by the span labeling approach \cite{spanviws} using XLM-RoBERTa \cite{conneau-etal-2020-unsupervised}, which is very slow when inference on CPU device. However, the well-known Vietnamese word segmenter toolkit \texttt{RDRsegmenter} achieved the F-score of 97.90\% by the rule approach with a fast speed for inference. In addition, from the above observations, word segmentation is used widely in many prior studies on Vietnamese sentiment classification, while some studies did not use word segmentation. Consequently, we have raised a research question, ``\textit{\textbf{Is word segmentation necessary for Vietnamese sentiment classification?}}''

To attempt to answer the question above, we used four Vietnamese word segmentation toolkits\footnote{Because of time limitations, we only selected the most recently published Vietnamese word segmentation toolkits. We will expand our research on other toolkits in future work.}. Firstly, we chose the fast and accurate Vietnamese word segmentation toolkit \texttt{RDRsegmenter} \cite{datnq1}. Secondly, we chose the Vietnamese word segmentation toolkit \texttt{uitnlp} \cite{vund2018}, which is proposed for ambiguity reduction and suffix capture. Lastly, we chose two well-known Vietnamese word segmentation toolkits, including \texttt{pyvi} and \texttt{underthesea} scientifically unpublished up to now. Regarding corpora, we chose two related to the Vietnamese sentiment classification problem, including the VLSP2016-SA corpus \cite{vlsp2016sa} of technical article reviews from the news and social media and the UIT-VSFC corpus of educational survey \cite{vsfc}. Regarding classifiers, we chose two traditional classifiers as baselines, including Naive Bayes (NB) and Support Vector Machines (SVMs) and the recent modern text encoder, namely, the Structured State Space Sequence model (S4) \cite{s4}.

In summary our contributions are the following:
\begin{itemize}
    \item Five pre-trained monolingual S4-based language models for Vietnamese, including one model without word segmentation, and four models using \texttt{RDRsegmenter} \cite{datnq1}, \texttt{uitnlp} \cite{vund2018}, \texttt{pyvi}, or \texttt{underthesea} toolkits in the pre-processing data phase.
    \item According to extensive experimental results on two corpora, including the VLSP2016-SA corpus of technical article reviews from the news and social media and the UIT-VSFC corpus of the educational survey, we have two suggestions:
    \begin{itemize}
        \item Using traditional classifiers like Naive Bayes or Support Vector Machines, word segmentation maybe not be necessary for the Vietnamese sentiment classification corpus, which comes from the social domain.
        \item Word segmentation is necessary for Vietnamese sentiment classification when word segmentation is considered pre-processing before using the BPE method and feeding into the deep learning model. By this way, the \texttt{RDRsegmenter} is the stable toolkit for word segmentation among the \texttt{uitnlp}, \texttt{pyvi}, and \texttt{underthesea} toolkits.
    \end{itemize}
\end{itemize}

\section{Pre-Trained Monolingual S4-Based Language Models For Vietnamese}
\label{sec:pre-trained}

In this paper, we focus on finding the effect of different word segmentation toolkits in the pre-processing data phase on deep learning models' Vietnamese sentiment classification performances. Hence, this section describes the models that applied different word segmentation toolkits, used the same S4-based language model architecture, and trained on the same pre-training data with the same optimization algorithm.

\subsection{Architecture}
\label{subsec:architecture}
This year, \citet{s4} suggested that the Structured State Space Sequence model (S4) has the potential to be an effective general sequence modeling solution. Hence, we chose S4 model \cite{s4} to develop pre-trained monolingual S4-based language models for Vietnamese. Besides, we have to train up to five monolingual S4-based language models for Vietnamese, including one model without word segmentation, and four models using \texttt{RDRsegmenter}, \texttt{uitnlp}, \texttt{pyvi}, or \texttt{underthesea} toolkits in the pre-processing data phase. Therefore, we used only two blocks of S4 layers (instead of 16 blocks in the original study \cite{s4}) alternated with position-wise feedforward layers, with a feature dimension of 256. Following \cite{s4}, we used a GLU activation after the S4 linear layer and used two S4 layers per block. The embedding and softmax layers were the Adaptive Embedding from \cite{DBLP:journals/corr/abs-1809-10853} with customized cutoffs 5000, 10000, 15000. The embedding size is 256, and the vocabulary size is 64K, as we will in the following subsection (\ref{subsec:pre-training-data}). Therefore, there are about 20M parameters for each such model, smaller than about six times compared with $\text{PhoBERT}_\text{base}$ \cite{phobert}.

\subsection{Pre-training Data}
\label{subsec:pre-training-data}
We used a 10 GB pre-training dataset, is the concatenation of two corpora, the first one is the Vietnamese Wikipedia corpus\footnote{\url{https://github.com/NTT123/viwik18/tree/viwik19}} ($\sim$ 1 GB), and the second corpus $(\sim$ 9 GB) is apart from the 18.6 GB Vietnamese news corpus\footnote{\url{https://github.com/binhvq/news-corpus\#full-txttitle--description--body-v1}}. To make a fair comparison of five pre-trained monolingual S4-based language models for Vietnamese mentioned in the previous subsection (\ref{subsec:architecture}), after doing pre-processing phase with word segmentation or without word segmentation, we apply the BPE method \cite{sennrich-etal-2016-neural} to segment the sentences from the pre-training dataset with subword units, using a vocabulary of 64K subword types.

\subsection{Optimization}

We inherit the implementation\footnote{\url{https://github.com/HazyResearch/state-spaces}} of S4-based language model from the study of \citet{s4}. We set a maximum length of 512 subword tokens when training and evaluating five pre-trained monolingual S4-based language models for Vietnamese. Following \citet{s4}, we optimize models using AdamW \cite{DBLP:journals/corr/abs-1711-05101} with a single cosine learning rate cycle with a maximum of 40 epochs and a number of warmup steps is 10000. The initial learning rate was set to 0.0005. We use a batch size of 128 with gradient accumulation steps of 16 on a V100 GPU (16 GB). The training is performed on the Google Colaboratory\footnote{\url{https://colab.research.google.com/}}. Lastly, we trained each S4-based language model for 10 days, except 14 days for the model without word segmentation in pre-processing phase.

\section{Experimental Setup}

We evaluate the performance with or without the word segmentation phase of two traditional classifiers (Naive Bayes and Support Vector Machines) and one deep learning classifier (S4-based language modeling), on two well-known Vietnamese sentiment corpora, including the VLSP2016-SA corpus of technical article reviews from the news and social media and the UIT-VSFC corpus of the educational survey. In this paper, we note again that we focus on finding the effect of applying or not applying the word segmentation phase in the pre-processing data phase on deep learning models' Vietnamese sentiment classification performances. This means we do not focus on finding new state-of-the-art results.

\subsection{Corpora}

\begin{table}[H]
\centering
\caption{Number Of Samples Of Each Class In VLSP2016-SA Corpus}
\label{tab:stat:vlsp2016}
\resizebox{0.85\columnwidth}{!}{%
\begin{tabular}{|l|r|r|r|r|}
\hline
 & \multicolumn{1}{l|}{Negative} & \multicolumn{1}{l|}{Neutral} & \multicolumn{1}{l|}{Positive} & \multicolumn{1}{l|}{Overall} \\ \hline
\textbf{Training} & 1,526 & 1,519 & 1,530 & 4,575 \\ \hline
\textbf{Validation} & 174 & 181 & 170 & 525 \\ \hline
\textbf{Test} & 350 & 350 & 350 & 1,050 \\ \hline
\end{tabular}%
}
\end{table}

Table~\ref{tab:stat:vlsp2016} presents the statistics of the VLSP2016-SA \cite{vlsp2016sa} corpus. The original corpus does not have a validation set. Therefore, we have split the original training set into the new training and validation sets.

\begin{table}[H]
\centering
\caption{Number Of Samples Of Each Class In UIT-VSFC Corpus}
\label{tab:stat:uitvsfc}
\resizebox{0.85\columnwidth}{!}{%
\begin{tabular}{|l|r|r|r|r|}
\hline
 & \multicolumn{1}{l|}{Negative} & \multicolumn{1}{l|}{Neutral} & \multicolumn{1}{l|}{Positive} & \multicolumn{1}{l|}{Overall} \\ \hline
\textbf{Training} & 5,325 & 458 & 5,643 & 11,426 \\ \hline
\textbf{Validation} & 705 & 73 & 805 & 1,583 \\ \hline
\textbf{Test} & 1,409 & 167 & 1,590 & 3,166 \\ \hline
\end{tabular}%
}
\end{table}

Table~\ref{tab:stat:uitvsfc} presents the statistics of the UIT-VSFC \cite{vsfc} corpus. We followed the splitting of training, validation, and test sets by \citet{vsfc}.

\subsection{Training}
For traditional classifiers, we used the combination of uni-grams, bi-grams, and tri-grams as the features for the classification problem. We used the \texttt{scikit-learn} library \cite{scikit-learn} for implementing\footnote{\url{https://scikit-learn.org/stable/}} the Naive Bayes and Support Vector Machines algorithms. For each corpus, we did not do word segmentation and did word segmentation using one of \texttt{RDRsegmenter}, \texttt{uitnlp}, \texttt{pyvi}, or \texttt{underthesea} toolkits. The experiment result is different for Support Vector Machines if we used different random states. Therefore, we report the average results of 100 runs with 100 random states for fair comparisons.

For modern classifiers, we fine-tuned five pre-trained monolingual S4-based language models for Vietnamese without word segmentation with appropriate word segmentation toolkits (\texttt{RDRsegmenter}, \texttt{uitnlp}, \texttt{pyvi}, or \texttt{underthesea}). We used the max-pooling vector of all contextual representations at the last layer of the S4-based language model for all positions of the input sentence as the feature vector for classification. We optimize models using AdamW \cite{DBLP:journals/corr/abs-1711-05101} with a single cosine learning rate cycle with a maximum of 5 epochs for the UIT-VSFC corpus and 10 epochs for the VLSP2016-SA corpus. The warmup proportion is 0.025. We set a maximum length of 8192 subwords for an input text and a batch size of 32 when fine-tuning both UIT-VSFC and VLSP2016-SA corpora. The initial learning rate is 0.001, and the weight decay of AdamW is 0.01. Notably, because we did not tune hyper-parameters in our experiments, we report the average results of 100 runs with 100 random states for fair comparisons. Finally, we have done fine-tuning experiments for 4 days.

\section{Experimental Results}

\subsection{Main Results}

\begin{table}[H]
\centering
\caption{Average Evaluation For Classification Models Over 100 Runs On Test Set Of VLSP2016-SA Corpus (\%)}
\label{tab:vlsp2016-sa-result}
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|l|ccc|}
\hline
\multicolumn{1}{|l|}{\multirow{2}{*}{\textbf{Model}}} & \multicolumn{1}{c|}{\multirow{2}{*}{\textbf{\begin{tabular}[c]{@{}c@{}}Word\\ Segmenter\end{tabular}}}} & \multicolumn{3}{c|}{\textbf{VLSP2016-SA}} \\ \cline{3-5} 
\multicolumn{1}{|l|}{} & \multicolumn{1}{c|}{} & \multicolumn{1}{c|}{\textbf{micro-$F_1$}} & \multicolumn{1}{c|}{\textbf{macro-$F_1$}} & \textbf{weighted-$F_1$} \\ \hline
\multirow{5}{*}{NB} & \texttt{None} & \multicolumn{1}{c|}{\textbf{67.24}} & \multicolumn{1}{c|}{\textbf{67.38}} & \textbf{67.38} \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{66.38} & \multicolumn{1}{c|}{66.39} & 66.39 \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{66.00} & \multicolumn{1}{c|}{66.00} & 66.00 \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{67.14} & \multicolumn{1}{c|}{67.13} & 67.13 \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{65.62} & \multicolumn{1}{c|}{65.62} & 65.62 \\ \hline
\multirow{5}{*}{SVMs} & \texttt{None} & \multicolumn{1}{c|}{\textbf{67.96}} & \multicolumn{1}{c|}{\textbf{67.99}} & \textbf{67.99} \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{67.56} & \multicolumn{1}{c|}{67.56} & 67.56 \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{67.33} & \multicolumn{1}{c|}{67.33} & 67.33 \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{67.93} & \multicolumn{1}{c|}{67.88} & 67.88 \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{66.77} & \multicolumn{1}{c|}{66.78} & 66.78 \\ \hline
\multirow{5}{*}{S4} & \texttt{None} & \multicolumn{1}{c|}{67.44} & \multicolumn{1}{c|}{67.42} & 67.42 \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{68.14} & \multicolumn{1}{c|}{68.07} & 68.07 \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{\textbf{68.51}} & \multicolumn{1}{c|}{\textbf{68.47}} & \textbf{68.47} \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{67.88} & \multicolumn{1}{c|}{67.83} & 67.83 \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{67.85} & \multicolumn{1}{c|}{67.85} & 67.85 \\ \hline
\end{tabular}
}
\end{table}

Table~\ref{tab:vlsp2016-sa-result} presents average evaluation on the VLSP2016-SA corpus. Firstly, the Naive Bayes classifier trained without the word segmentation phase achieved higher performance than other Naive Bayes classifiers trained with the word segmentation phase. This also happened for the Support Vector Machines classifier. Secondly, S4-based classifiers trained with the word segmentation phase achieved higher performance than the S4-based classifier trained without the word segmentation phase. Especially, the S4-based classifier with the \texttt{uitnlp} toolkit achieved the higher result compared with \texttt{RDRsegmenter}, \texttt{pyvi}, and \texttt{underthesea}.

\begin{table}[H]
\centering
\caption{Average Evaluation For Classification Models Over 100 Runs On Test Set Of UIT-VSFC Corpus (\%)}
\label{tab:uit-vsfc-result}
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|l|ccc|}
\hline
\multicolumn{1}{|l|}{\multirow{2}{*}{\textbf{Model}}} & \multicolumn{1}{c|}{\multirow{2}{*}{\textbf{\begin{tabular}[c]{@{}c@{}}Word\\ Segmenter\end{tabular}}}} & \multicolumn{3}{c|}{\textbf{UIT-VSFC}} \\ \cline{3-5} 
\multicolumn{1}{|l|}{} & \multicolumn{1}{c|}{} & \multicolumn{1}{c|}{\textbf{micro-$F_1$}} & \multicolumn{1}{c|}{\textbf{macro-$F_1$}} & \textbf{weighted-$F_1$} \\ \hline
\multirow{5}{*}{NB} & \texttt{None} & \multicolumn{1}{c|}{86.26} & \multicolumn{1}{c|}{59.83} & 84.07 \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{86.99} & \multicolumn{1}{c|}{60.32} & 84.77 \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{87.11} & \multicolumn{1}{c|}{60.79} & 84.95 \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{\textbf{87.18}} & \multicolumn{1}{c|}{\textbf{60.83}} & \textbf{84.99} \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{86.99} & \multicolumn{1}{c|}{60.33} & 84.78 \\ \hline
\multirow{5}{*}{SVMs} & \texttt{None} & \multicolumn{1}{c|}{89.17} & \multicolumn{1}{c|}{73.48} & 88.55 \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{\textbf{89.36}} & \multicolumn{1}{c|}{72.96} & 88.60 \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{89.35} & \multicolumn{1}{c|}{73.05} & \textbf{88.62} \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{89.29} & \multicolumn{1}{c|}{\textbf{73.10}} & 88.54 \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{89.27} & \multicolumn{1}{c|}{72.58} & 88.47 \\ \hline
\multirow{5}{*}{S4} & \texttt{None} & \multicolumn{1}{c|}{90.88} & \multicolumn{1}{c|}{75.82} & 90.11 \\ \cline{2-5} 
 & \texttt{RDRsegmenter} & \multicolumn{1}{c|}{\textbf{91.62}} & \multicolumn{1}{c|}{\textbf{77.84}} & \textbf{90.99} \\ \cline{2-5} 
 & \texttt{uitnlp} & \multicolumn{1}{c|}{91.36} & \multicolumn{1}{c|}{77.17} & 90.68 \\ \cline{2-5} 
 & \texttt{pyvi} & \multicolumn{1}{c|}{91.40} & \multicolumn{1}{c|}{77.74} & 90.80 \\ \cline{2-5} 
 & \texttt{underthesea} & \multicolumn{1}{c|}{91.35} & \multicolumn{1}{c|}{77.26} & 90.72 \\ \hline
\end{tabular}
}
\end{table}

Table~\ref{tab:uit-vsfc-result} presents average evaluation on the UIT-VSFC corpus. As we can see in Table~\ref{tab:uit-vsfc-result}, the Naive Bayes, Suppor Vector Machine, and S4-based classifier with the word segmentation phase achieved higher performance than other classifiers without the word segmentation phase. Especially, the S4-based classifier with the \texttt{RDRsegmenter} toolkit achieved the higher result compared with \texttt{uitnlp}, \texttt{pyvi}, and \texttt{underthesea}.

\subsection{Discussion}

As we can see in Table~\ref{tab:vlsp2016-sa-result}, using word segmentation for traditional classifiers like Naive Bayes and Support Vector Machines maybe achieve a lower performance than classifiers without word segmentation for the Vietnamese sentiment classification task. According to the study of \citet{vund2018}, the \texttt{uitnlp} achieved the higher F-score of word segmentation than \texttt{RDRsegmenter} \cite{datnq1} on the Vietnamese Treebank corpus for word segmentation \cite{nguyen-etal-2009-building}. However, the VLSP2016-SA corpus of technical article reviews from the news and social media. Hence, the social corpus's word segmentation criteria may differ from the Vietnamese Treebank corpus for word segmentation \cite{nguyen-etal-2009-building}. Finally, the potential reason explaining the traditional classifier with \texttt{uitnlp} toolkit achieved lower performance than the classifier with \texttt{RDR} is that the \texttt{uitnlp} toolkit was trained with over-fitting on the Vietnamese Treebank corpus for word segmentation \cite{nguyen-etal-2009-building}, which is not good when \texttt{uitnlp} toolkit faces the social corpus. This suggests that building Vietnamese word segmentation and part-of-speech tagging for social media text, like the study of \citet{10.1007/978-3-319-46675-0_26} is worthy of respect and attention.

In such an analytical way, in Table~\ref{tab:uit-vsfc-result}, the Naive Bayes, Support Vector Machine, and S4-based classifier with the word segmentation phase achieved higher performance than other classifiers without the word segmentation phase. This can be explained by the word segmentation criteria in the educational survey domain being close to the Vietnamese Treebank corpus for word segmentation \cite{nguyen-etal-2009-building}. Because the UIT-VSFC corpus was normalized by humans \cite{vsfc}.

\section{Conclusion}

This paper attempts to answer whether word segmentation is necessary for Vietnamese sentiment classification. According to comprehensive experimental results on two corpora, including the VLSP2016-SA corpus of technical article reviews from the news and social media and the UIT-VSFC corpus of the educational survey, we have two suggestions. Firstly, using traditional classifiers like Naive Bayes or Support Vector Machines, word segmentation maybe not be necessary for the Vietnamese sentiment classification corpus, which comes from the social domain. Secondly, word segmentation is necessary for Vietnamese sentiment classification when word segmentation is used before using the BPE method and feeding into the deep learning model. In this way, the \texttt{RDRsegmenter} is the stable toolkit for word segmentation among the \texttt{uitnlp}, \texttt{pyvi}, and \texttt{underthesea} toolkits because the \texttt{RDRsegmenter} achieved the stable performance on sentiment classification task in our experiments and has fast inference speed \cite{datnq1}.

\section*{Acknowledgement}
This research was supported by The VNUHCM-University of Information Technology's Scientific Research Support Fund.

{
    \balance
    \footnotesize
    \bibliography{main.bib}
}

\end{document}